Shift Summary for Day shift, Sept 1, 1999

S. Elizabeth Turner (eturner@kudzu)
Wed, 1 Sep 1999 15:36:12 -0500

Shift Summary for Day shift, Sept 1, 1999

On Shift: SET, AE

Hours of beam available at all intensities: 1 hr
Hours of beam available at nominal intensity: 1 hr
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 1 hr

Accounting of lost time:
No beam for most of shift due to TeV studies.
Muon runs interrupted by DA problems, laser scans
interrupted by pipeline errors due to lack of spill
signal after beam went down.
========================================================================
Runs ended this shift:
14391 MU_4E12
14392 MU_4E12
14393 LASER_SCAN (aborted)
14394 LASER_SCAN (aborted)
many junk debugging runs

Tapes brought to FCC this shift:
(including tapes for online split)
QKS372-374
QKB129
========================================================================
Accesses during shift:
Albert & Bud to check for lamps.
Elizabeth and Rick to look at latches.
Ed and Julie to swap divider.
========================================================================
Overview of shift:
Scheduled downtime, some successful debugging of various
problems, testing DA configs for special runs.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14391 7:20 stop run due to DA problems

14392 start new muon run
7:35 cycle power on repeater to fix DA probs
7:58 begin scheduled downtime, end muon run

14393 laser scan, had to be aborted b/c of
pipeline errors -- errors were due to
lack of spill signal

14394 ditto.

8:50 MCR says no internal clock today b/c of
TeV studies so no spill signal all day

9:55 RSK and SET make access -- put HA 29 back
the way it was, look at latches to help
debug crate 8

10:20 crate 8 (trigger latches) is fixed --
problem was that CTI was terminating L2
signal before it reached latches

10:45 unloaded tapes. only QKS and QKB tapes
will be taken to FCC as others are empty

10:45 tried to init special run configurations
as a test, eventually both BEAM_MU_LONG
and BEAM_MAGOFF work

11:30 Changed HV for CSI450 from 1700 to 1546
Swapped divider for CSI450, HV & Q/E
constants restored to values from run 14350.
========================================================================
Follow up Notes on Previous Shift

laser scan -- we were unable to get a laser scan -- if spill signal
returns before beam, try to get this in.

========================================================================
Items to Follow up on Next Shift

Come up in BEAM_MAGOFF config. Page Rick K. at the beginning of
this run.